OpenStudy (anonymous):

missing coordinate.......2x+y=1 (0,y)

OpenStudy (anonymous):

Substitute 0 in for x. 2*0 + y = 1 What does y equal then?

OpenStudy (anonymous):

Same as before, plug 0 in for x, what does y equal.

OpenStudy (anonymous):

0+y=1

OpenStudy (anonymous):

Right, so if y + 0 = 1, what must y equal?

OpenStudy (anonymous):

1